OpenStudy (anonymous):

In her wallet, Ms. Thompson has one-dollar, five-dollar, and ten-dollar bills, totally $171. She has the same number of five-dollar bills as one-dollar and ten-dollar bills put together. If she has 30 bills in all, how many bills of each denomination does she have?

OpenStudy (anonymous):

i got 6 - 1 dollar bills 15 fives dollar bill and 9 ten dollars

sammixboo (sammixboo):

Let's check. 6 one-dollars = $6 15 five-dollars = $75 9 ten-dollars = $90 Add them all up to get $171, so that is correct. Add the number of one-dollar bills and the number of ten-dollar bills together. 6 + 9 = 15, which is the number of five-dollar bills, so that is correct as well. Add all the numbers of bills together, 6 + 9 + 15 = 15 + 15 = 30, so that is correct So yeah!

sammixboo (sammixboo):

Oh and... \(\huge \rlap{\bbox[2.9pt, ,border:6.8px solid #2aacd7 ]{\Huge\color{blueviolet}{\rm \fbox{Welcome}}}}\hspace{127pt}\huge \rlap{\bbox[3.3pt, ,border:7px solid #2aacd7 ]{\Huge\color{blueviolet}{\rm \fbox{to}}}}\) \(\huge \color{blueviolet}{\heartsuit^{\LARGE \heartsuit^{\Large \heartsuit}}}\) \(\huge \color{blueviolet}{\heartsuit^{\LARGE \heartsuit^{\Large \heartsuit}}}\) \(\huge \huge \rlap{\bbox[5pt, ,border:8px solid #2aacd7 ]{\Huge\color{#84c73c}{\rm \fbox{OpenStudy}}}}\)